Two cylindrical flower pots need to be painted. The flower pots are 1.5 ft in diameter and 2.5 ft high. A one litre can of paint covers 100 ft2. Will one can be enough to paint the two flower pots with two coats of paint?
1
Expert's answer
2011-12-09T09:54:25-0500
Let's make the following denotations:

D = 1.5 ft
H = 2.5 ft
S = 100 ft²

D = 2*R ==> R = 1.5/2 = 0.75 ft - the radius of a pot.

The surface area of a pot is

A = 2*pi*R^2 + 2*pi*R*H = 2*pi*0.75^2 + 2*pi*0.75*2.5 ≈ 15.3154 ft².

So, two flower pots with two coats of paint need

V = (A*2*2)/S = 15.3154*4/100 = 0.6126 litres of paint.

So, one litre of paint is enough to paint the two flower pots with two coats of paint.